# Analysis of Linear Mode Connectivity via Permutation-Based Weight Matching: With Insights into Other Permutation Search Methods

### v4

@@ 4 Analysis of WM
The previous section demonstrates that the establishment of LMC by WM is not due to the reduction in $L_{2}$ distance itself, but rather because WM helps find permutations that result in a smaller barrier between the two models. To better understand why WM reduces the barrier, we first analyze WM by performing SVD on the weights of each layer of the model in Section 4.1. Then, in Section 4.2, we show that the singular value distribution of each layer are almost identical across independently trained models, and that the primary differences between the models are due to variations in their singular vectors. In Section 4.3, we demonstrate that WM preferentially aligns the directions of singular vectors with large singular values between the weights of the two models. Finally, in Section 4.4, we explain that aligning singular vectors with large singular values makes LMC more achievable because these singular vectors predominantly influence the outputs of the hidden layers of the models.
@@ 4.1 Analysis Based on SVD
The basic idea in analyzing WM is to perform SVD on the weight in each layer. Although using SVD for analysis might seem overly simplistic given that WM reduces the $L_{2}$ distance, this approach provides important insights that are explored in subsequent sections. In WM, permutation matrices are searched to minimize Equation 2. We denote the SVDs of $\bm{W}_{\ell}^{(a)}$ and $\bm{W}_{\ell}^{(b)}$ by $\bm{W}_{\ell}^{(a)}=\bm{U}_{\ell}^{(a)}\bm{S}_{\ell}^{(a)}(\bm{V}_{\ell}^{(a)})^{\top}=\sum_{i}\bm{u}_{\ell,i}^{(a)}s_{\ell,i}^{(a)}(\bm{v}_{\ell,i}^{(a)})^{\top}$ and $\bm{W}_{\ell}^{(b)}=\bm{U}_{\ell}^{(b)}\bm{S}_{\ell}^{(b)}(\bm{V}_{\ell}^{(b)})^{\top}=\sum_{j}\bm{u}_{\ell,j}^{(b)}s_{\ell,j}^{(b)}(\bm{v}_{\ell,j}^{(b)})^{\top}$, respectively. Here, we assume that the singular values are ordered in descending order (i.e., for all $\ell\in[L]$, $s_{\ell,1}^{(a)}\geq s_{\ell,2}^{(a)}\geq\dots\geq s_{\ell,n}^{(a)}$ and $s_{\ell,1}^{(b)}\geq s_{\ell,2}^{(b)}\geq\dots\geq s_{\ell,n}^{(b)}$, where $n$ is the number of singular values). Then, we have
@@
Equation 4 shows that the permutation matrices $\bm{P}_{\ell}$ and $\bm{P}_{\ell-1}$ are multiplied by the left and right singular vectors of the model $\bm{\theta}_{b}$, respectively. If the singular values of both models are almost equal (i.e., for all $\ell$ and $i$, $s_{\ell,i}^{(a)}=s_{\ell,i}^{(b)}$) and we can find permutation matrices such that the left and right singular vectors of the two models match, then the two models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ will be exactly the same and LMC clearly holds. Therefore, in the following, we will discuss the differences in (1) singular values and (2) singular vectors of independently trained models.
@@ 4.2 Differences Between Singular Values of Two Models
First, we investigate the differences between the singular values of two independently trained models. To this end, ten models are trained independently under identical conditions except for the seed, and their singular values are compared. Figure 1 plots the singular values in the second layer of independently trained models in descending order. The evaluation results for all the layers are shown in Figure 7. As can be seen in the figures, in the hidden layers, the singular values are very close across all models. Therefore, the differences in singular values between the models are not a major issue in achieving LMC.
@@ 4.3 Singular-Vector Alignment
In the previous subsection, we confirmed that the distributions of the singular values of the weights of the two independently trained models were almost equal. In this subsection, we will show that the permutations found by WM only align the large singular vectors of the two models. We will also show experimentally that WM cannot align all singular vectors between the two models, and therefore cannot reduce the $L_{2}$ distance to zero.
@@ 4.4 Importance of Singular Vectors in LMC
In the previous section, we mentioned that WM aligns the directions of singular vectors with large singular values. This section clarifies why these singular vectors, rather than $L_{2}$ distances, play a crucial role in establishing LMC.
@@
Thus, as long as the directions of the $k$-th singular vectors of the two weights are aligned (i.e., $\bm{v}_{\ell,k}^{(a)}=\bm{v}_{\ell,k}^{(b)}$ and $\bm{u}_{\ell,k}^{(a)}=\bm{u}_{\ell,k}^{(b)}$ hold), the outputs of the two layers will coincide regardless of the other singular vectors. Note that the $L_{2}$ distance between the two weights is not necessarily close to zero since the directions of the other singular vectors need not be aligned. In fact, in Appendix C, we provide an example where the $L_{2}$ distance is not close to zero even though the output is zero.
@@
In the context of WM, it is desirable that the input vector has the same direction as the right singular vectors with large singular values because WM preferentially aligns these singular vectors of the two weights. To verify this, we experimentally investigate the relationship between the directions of the right singular vectors and that of the input vector. Figure 4 shows the value of $\mathbb{E}(\bm{v}_{\ell,i}^{\top}\bm{z}_{\ell-1})^{2}$ for the $i$-th right singular vector $\bm{v}_{\ell,i}$ in the second layer (i.e., $\ell=2$) and the corresponding hidden layer input $\bm{z}_{\ell-1}$ for each model. The results show that in the hidden layer, the singular vectors with large singular values have large inner products with the input vectors, which indicates that the permutations found by WM make LMC more feasible. In particular, Figure 3 shows that by aligning the directions of the singular vectors between the two models $\bm{\theta_{a}}$ and $\bm{\theta_{b}}$ using WM, the directions of the singular vectors with large singular values of the models before and after merging (e.g., $\bm{\theta}_{a}$ and $(\bm{\theta}_{a}+\bm{\theta}_{b})/2$) are well aligned. This suggests that the hidden layer outputs of the models before and after merging are closer, which contributes to the establishment of LMC.
Some studies (Ainsworth et al., 2023; Entezari et al., 2022) have empirically observed that increasing the model width makes it easier to satisfy LMC using WM. In Section H.3, we provide an analysis to explain this observation in terms of singular-vector alignment. Furthermore, Qu & Horvath (2024) showed that strengthening weight decay and increasing the learning rate make it easier for LMC to be established through WM. Section H.4 demonstrates that increasing these values reduces the proportion of large singular values in the weights of each layer, facilitating the alignment of the corresponding singular vectors through WM, and thus making LMC easier to achieve.
@@ 5 Activation Matching
Activation Matching (AM) searches for a permutation $\pi^{*}$ based on the following equation:
@@
The minimization of Equation 10 is closely related to Theorem 4.2. Specifically, in a permutation search for the $\ell$-th layer, if we assume that the outputs $\bm{z}_{\ell-1}^{(a)}$ and $\bm{z}_{\ell-1}^{(b)}$ from the previous layer are sufficiently close under the permutation $\bm{P}_{\ell-1}$ (i.e., $\bm{z}_{\ell-1}^{(a)}\approx\bm{P}_{\ell-1}\bm{z}_{\ell-1}^{(b)}$), then minimizing the right-hand side of Equation 9 becomes equivalent to reducing the objective function in Equation 10. In other words, similar to WM, AM may search for permutations that align the singular vectors with large singular values between two models. To verify this, the results of model merging using AM are presented in Table 3. The experimental settings are the same as those used for WM in Section 3.2. Additionally, to evaluate how well the singular vectors align through permutation, the $R$ calculation results are shown in Figures 14 and 15. These results closely resemble those for WM in Tables 1, 2 and 3, suggesting that the reason AM achieves LMC is likely the same as for WM.
@@ 6 Comparison with Straight-Through Estimator
Dataset Network Barrier ($\lambda=1/2$) $L_{2}$ dist. w/o STE $L_{2}$ dist. w/ STE $R(\bm{\theta}_{a},\pi(\bm{\theta}_{b}))$ ($\gamma=0.3)$
@@
This section discusses the relationship between the straight-through estimator (STE), a more direct method for finding permutations, and WM in terms of singular vectors. STE uses a dataset to find permutation matrices with a small barrier value. This section also explains that STE and WM are based on fundamentally different principles, and shows how this difference significantly impacts LMC among three or more models.
@@ 6.1 Straight-through Estimator (STE)
Since Equation 11 is difficult to solve directly, Ainsworth et al. (2023) proposed a method to approximate the solution. Later, Peña et al. (2023) proposed a method to solve Equation 11 directly using Sinkhorn’s algorithm. We adopt the latter method based on Sinkhorn’s algorithm, which we refer to as STE in this paper.
@@ 6.2 Experimental Results of Model Merging by STE
(a) Results of WM.
(b) Results of STE.
The experimental results of model merging using STE are shown in Table 2. This table also shows the $L_{2}$ distance and the $R$ value between the two models before and after the permutation. Despite the relatively small barrier value, Table 2 shows that the $L_{2}$ distance between the two models before and after permutation hardly changes compared to the results with WM shown in Table 1. Since $R(\bm{\theta}_{a},\pi(\bm{\theta}_{b}))$ is nearly zero, the singular vectors between the two models are likely not aligned at all. Therefore, the reason for satisfying LMC by STE is completely different from that of WM.
@@ 6.3 LMC among Three Models
We performed model merging experiments among three models to confirm the validity of the above discussion. First, Figure 5 presents the results of examining how well the singular vectors are aligned in each model pair by WM. Since the models $\bm{\theta}_{b}$ and $\bm{\theta}_{c}$ are matched to $\bm{\theta}_{a}$ through WM, it is expected that $R(\bm{\theta}_{a},\pi_{b}(\bm{\theta}_{b}))$ and $R(\bm{\theta}_{a},\pi_{c}(\bm{\theta}_{c}))$ would be large. On the other hand, although $\bm{\theta}_{b}$ and $\bm{\theta}_{c}$ were not explicitly aligned, $R(\pi_{b}(\bm{\theta}_{b}),\pi_{c}(\bm{\theta}_{c}))$ is clearly greater than zero, indicating that the directions of these two singular vectors are indirectly aligned by WM. From this result, the barrier between the models $\pi_{b}(\bm{\theta}_{b})$ and $\pi_{c}(\bm{\theta}_{c})$ is expected to be small. To confirm this, Figure 6 shows the test accuracy landscape around $\bm{\theta}_{a}$, $\pi_{b}(\bm{\theta}_{b})$, and $\pi_{c}(\bm{\theta}_{c})$. The results of model merging at $\lambda=1/2$ for each pair of models are also shown in Table 4. Table 4 shows the mean and standard deviation of model merging three times. As can be seen from Figure 6, the barrier between $\pi_{b}(\bm{\theta}_{b})$ and $\pi_{c}(\bm{\theta}_{c})$ is smaller with WM than with STE. This means that there is a significant difference between the principles of permutations obtained by WM and STE. Figure 6 also shows that the landscape of test accuracy is flatter around the three models with WM than with STE. Therefore, WM is likely to be more advantageous, especially for merging three or more models.
@@ 7 Conclusion
This paper analyzed why linear mode connectivity (LMC) is satisfied through permutation search with weight matching (WM). First, we demonstrated that WM does not reduce the distance between the weights of two models as significantly as previously thought. We then analyzed WM using singular value decomposition (SVD) and found that WM aligns the directions of singular vectors with large singular values, which plays a crucial role in achieving LMC. Additionally, we showed that the reason LMC is established in AM is the same as for WM. Finally, we discussed the difference between STE and WM from the perspective of singular vectors.
@@ Appendix A Extended Related Work
Several studies (Garipov et al., 2018; Draxler et al., 2018; Freeman & Bruna, 2017) have found that different neural network solutions can be connected by nonlinear paths with almost no increase in loss. Nagarajan & Kolter (2019) first discovered that solutions can be connected by linear paths with an almost constant loss value when training models on MNIST with the same random initial values. Later, Frankle et al. (2020) demonstrated experimentally that LMC is not always satisfied between two SGD solutions, even with the same initial parameters, depending on the datasets and model architectures. However, they also showed that if a single model is trained for a certain period and then two models are trained independently from this pre-trained model as a starting point, they are linearly mode-connected. Furthermore, Frankle et al. (2020) explored the relationship between LMC and the lottery-ticket hypothesis (Frankle & Carbin, 2019). Entezari et al. (2022) conjectured that LMC is satisfied with a high probability between two SGD solutions by accounting for permutation symmetries in the hidden layers. Subsequently, Ainsworth et al. (2023) proposed a WM method by formulating neuron alignment as a bipartite graph matching problem and solving it approximately. Later, Peña et al. (2023) suggested using Sinkhorn’s algorithm to solve the WM directly. Some previous studies (Ainsworth et al., 2023; Crisostomi et al., 2024) have also proposed permutation search methods for achieving LMC between multiple models. However, all of these methods use WM, and no methods have been proposed that use information from the loss function, such as STE. Therefore, in this paper, we created a pair of models and performed a permutation search to clarify the differences between WM and STE in a fair manner. The investigation of permutation search methods for multiple models is a future work.
@@
While several papers (Venturi et al., 2019; Nguyen et al., 2019; Nguyen, 2019; Kuditipudi et al., 2019) have discussed nonlinear mode connectivity, there is little theoretical analysis on LMC. Ferbach et al. (2024) provided an upper bound on the minimal width of the hidden layer to satisfy LMC. However, to prove this, they assumed the independence of all neuron’s weight vectors inside a given layer. It is unlikely that this assumption holds for models after training. Zhou et al. (2023) introduced the concept of layerwise linear feature connectivity (LLFC) and showed that LLFC implies LMC. Additionally, Zhou et al. (2023) demonstrated that if weak additivity for ReLU activation and the commutativity property are satisfied, then both LLFC and LMC are satisfied. However, we show that the $L_{2}$ distance between the models after permutation is not close enough to satisfy the commutativity property. This motivated us to investigate the relationship between LMC and WM.
@@
Empirically, some previous studies (Tukan et al., 2020; Arora et al., 2018; Alvarez & Salzmann, 2017; Yu et al., 2017; Denton et al., 2014) on the compression of trained DNN models have pointed out that even when the weights are replaced with low-rank matrices, the accuracy does not decrease significantly. This suggests that SGD has an implicit bias toward reducing the model weights to low-rank. Galanti et al. (2023) and Timor et al. (2023) further discuss the low-rank effect on the weights of trained models introduced by weight decay and a small initialization scale. In particular, Galanti et al. (2023) state that the weights become more low-rank by strengthening the weight decay and increasing the learning rate, and this is expected to help establish LMC via WM (which we experimentally confirm in Section H.4).

### v5

@@ 4 Analysis of WM
The previous section demonstrates that the establishment of LMC by WM is not due to the reduction in $L^{2}$ distance itself, but rather because WM helps find permutations that result in a smaller barrier between the two models. To better understand why WM reduces the barrier, we first analyze WM by performing SVD on the weights of each layer of the model in Section 4.1. Then, in Section 4.2, we show that the singular value distribution of each layer is almost identical across independently trained models, and that the primary differences between the models are due to variations in their singular vectors. In Section 4.3, we demonstrate that WM preferentially aligns the directions of singular vectors with large singular values between the weights of the two models. Finally, in Section 4.4, we explain that aligning singular vectors with large singular values makes LMC more achievable because these singular vectors predominantly influence the outputs of the hidden layers of the models.
@@ 4.1 Analysis Based on SVD
The basic idea in analyzing WM is to perform SVD on the weight in each layer. Although using SVD for analysis might seem overly simplistic given that WM reduces the $L^{2}$ distance, this approach provides important insights that are explored in subsequent sections. In WM, permutation matrices are searched to minimize Equation 2. We denote the SVDs of $\bm{W}_{\ell}^{(a)}$ and $\bm{W}_{\ell}^{(b)}$ by $\bm{W}_{\ell}^{(a)}=\bm{U}_{\ell}^{(a)}\bm{S}_{\ell}^{(a)}(\bm{V}_{\ell}^{(a)})^{\top}=\sum_{i}\bm{u}_{\ell,i}^{(a)}s_{\ell,i}^{(a)}(\bm{v}_{\ell,i}^{(a)})^{\top}$ and $\bm{W}_{\ell}^{(b)}=\bm{U}_{\ell}^{(b)}\bm{S}_{\ell}^{(b)}(\bm{V}_{\ell}^{(b)})^{\top}=\sum_{j}\bm{u}_{\ell,j}^{(b)}s_{\ell,j}^{(b)}(\bm{v}_{\ell,j}^{(b)})^{\top}$, respectively. Here, we assume that the singular values are ordered in descending order (i.e., for all $\ell\in[L]$, $s_{\ell,1}^{(a)}\geq s_{\ell,2}^{(a)}\geq\dots\geq s_{\ell,n}^{(a)}$ and $s_{\ell,1}^{(b)}\geq s_{\ell,2}^{(b)}\geq\dots\geq s_{\ell,n}^{(b)}$, where $n$ is the number of singular values). Then, we have
@@
Equation 4 shows that the permutation matrices $\bm{P}_{\ell}$ and $\bm{P}_{\ell-1}$ are multiplied by the left and right singular vectors of the model $\bm{\theta}_{b}$, respectively. The $L^{2}$ distance between the models is expressed by the difference in singular values and singular vectors between the two models, as indicated by Equation 4. Therefore, in the following, we will discuss the differences in (1) singular values and (2) singular vectors of independently trained models.
@@ 4.2 Differences Between Singular Values of Two Models
First, we investigate the differences between the singular values of two independently trained models. To this end, ten models are trained independently under identical conditions except for the seed, and their singular values are compared. Figure 1 plots the singular values in the second layer of independently trained models in descending order. The evaluation results for all the layers are shown in Figure 7. As can be seen in the figures, in the hidden layers, the singular values are very close across all models. Therefore, the differences in singular values between the models are not a significant obstacle to reducing the distance between the two models to zero.
@@ 4.3 Singular-Vector Alignment
In the previous subsection, we confirmed that the distributions of the singular values of the weights of the two independently trained models were almost equal. In this subsection, we will show that the permutations found by WM preferentially align the dominant singular vectors of the two models, and cannot align all singular vectors between the two models. Therefore, WM cannot reduce the $L^{2}$ distance to zero.
@@ 4.4 Importance of Singular Vectors in LMC
In the previous section, we mentioned that WM aligns the directions of singular vectors with large singular values. This section clarifies why these singular vectors, rather than $L^{2}$ distances, play a crucial role in establishing LMC.
@@
Thus, as long as the directions of the $k$-th singular vectors of the two weights are aligned (i.e., $\bm{v}_{\ell,k}^{(a)}=\bm{v}_{\ell,k}^{(b)}$ and $\bm{u}_{\ell,k}^{(a)}=\bm{u}_{\ell,k}^{(b)}$ hold), the outputs of the two layers will coincide regardless of the other singular vectors. Note that the $L^{2}$ distance between the two weights is not necessarily close to zero since the directions of the other singular vectors need not be aligned. In fact, in Appendix C, we provide an example where the $L^{2}$ distance is not close to zero even though the output is zero.
@@
In the context of WM, it is desirable that the input vector has the same direction as the right singular vectors with large singular values because WM preferentially aligns these singular vectors of the two weights. To verify this, we experimentally investigate the relationship between the directions of the right singular vectors and that of the input vector. Figure 4 shows the value of $\mathbb{E}(\bm{v}_{\ell,i}^{\top}\bm{z}_{\ell-1})^{2}$ for the $i$-th right singular vector $\bm{v}_{\ell,i}$ in the second layer (i.e., $\ell=2$) and the corresponding hidden layer input $\bm{z}_{\ell-1}$ for each model. The results show that in the hidden layer, the singular vectors with large singular values have large inner products with the input vectors, which indicates that the permutations found by WM make LMC more feasible. [^3] As shown in Figures 11 and 12, this tendency only occurs when the model is sufficiently wide. This suggests that LMC is unlikely to be established in WM unless the model is wide enough. In particular, Figure 3 shows that by aligning the directions of the singular vectors between the two models $\bm{\theta_{a}}$ and $\bm{\theta_{b}}$ using WM, the directions of the singular vectors with large singular values of the models before and after merging (e.g., $\bm{\theta}_{a}$ and $(\bm{\theta}_{a}+\bm{\theta}_{b})/2$) are well aligned. This suggests that the hidden layer outputs of the models before and after merging are closer, which contributes to the establishment of LMC.
Some studies (Ainsworth et al., 2023; Entezari et al., 2022) have observed that increasing the model width makes it easier to satisfy LMC using WM. In Section H.4, we provide an empirical analysis to explain this observation in terms of singular-vector alignment. Furthermore, Qu & Horvath (2024) showed that strengthening weight decay and increasing the learning rate makes it easier for LMC to be established through WM. Section H.5 demonstrates empirically that increasing these values reduces the proportion of large singular values in the weights of each layer, facilitating the alignment of the corresponding singular vectors through WM, and thus making LMC easier to achieve.
@@ 5 Activation Matching
AM searches for a permutation $\pi^{*}$ based on the following equation:
@@
The minimization of Equation 10 is related to Theorem 4.2. Specifically, in a permutation search for the $\ell$-th layer, if we assume that the outputs $\bm{z}_{\ell-1}^{(a)}$ and $\bm{z}_{\ell-1}^{(b)}$ from the previous layer are sufficiently close under the permutation $\bm{P}_{\ell-1}$ (i.e., $\bm{z}_{\ell-1}^{(a)}\approx\bm{P}_{\ell-1}\bm{z}_{\ell-1}^{(b)}$), then minimizing the right-hand side of Equation 9 becomes equivalent to reducing the objective function in Equation 10. In other words, similar to WM, AM may search for permutations that align the singular vectors with large singular values between two models. To verify this, the results of model merging using AM are presented in Table 4. The experimental settings are the same as those used for WM in Section 3.2. Additionally, to evaluate how well the singular vectors align through permutation, the $R$ calculation results are shown in Figures 15 and 16. These results closely resemble those for WM in Tables 1, 2 and 3, suggesting that the reason AM achieves LMC is likely to be similar to that for WM.
@@ 6 Comparison with Straight-Through Estimator
Dataset Network Barrier ($\lambda=1/2$) $L^{2}$ dist. w/o STE $L^{2}$ dist. w/ STE $R(\bm{\theta}_{a},\pi(\bm{\theta}_{b}))$ ($\gamma=0.3)$
@@
This section discusses the relationship between the straight-through estimator (STE), a more direct permutation search method, and WM in terms of singular vectors. STE uses a dataset to find permutations with a small barrier value. We also explain that STE and WM are based on fundamentally different principles and show how this difference impacts LMC among three or more models.
@@ 6.1 Straight-through Estimator (STE)
Since Equation 11 is difficult to solve directly, Ainsworth et al. (2023) proposed a method to approximate the solution. Later, Guerrero-Peña et al. (2023) proposed a method to solve Equation 11 directly using Sinkhorn’s algorithm. We adopt the latter method, which we refer to as STE in this paper.
@@ 6.2 Experimental Results of Model Merging by STE
Table 3: Loss and accuracy barriers between $\pi_{b}(\bm{\theta}_{b})$ and $\pi_{c}(\bm{\theta}_{c})$. The table shows the mean and standard deviation over three model merging trials.
Loss barrier ($(\pi_{b}(\bm{\theta}_{b})+\pi_{c}(\bm{\theta}_{c}))/2$) Accuracy barrier ($(\pi_{b}(\bm{\theta}_{b})+\pi_{c}(\bm{\theta}_{c}))/2$)
Dataset Network WM STE WM STE
CIFAR10 VGG11 $0.141\pm 0.141$ $2.172\pm 0.989$ $10.12\pm 5.117$ $32.013\pm 8.193$
ResNet20 $0.294\pm 0.098$ $1.693\pm 0.168$ $7.23\pm 0.99$ $34.483\pm 2.426$
FMNIST MLP $-0.174\pm 0.051$ $0.023\pm 0.118$ $4.337\pm 1.434$ $15.97\pm 1.724$
MNIST MLP $-0.031\pm 0.003$ $0.017\pm 0.014$ $0.475\pm 0.069$ $2.312\pm 0.457$
The experimental results of model merging using STE are shown in Table 2. This table also shows the $L^{2}$ distance and the $R$ value between the two models before and after the permutation. Despite the relatively small barrier value, Table 2 shows that the $L^{2}$ distance between the two models before and after permutation hardly changes compared to the results with WM shown in Table 1. Since $R(\bm{\theta}_{a},\pi(\bm{\theta}_{b}))$ is nearly zero, the singular vectors between the two models are likely not aligned at all. Therefore, the reason for satisfying LMC by STE is completely different from that of WM.
@@ 6.3 LMC among Three Models
We performed model merging experiments among three models to confirm the validity of the above discussion. First, Figure 5 presents the results of examining how well the singular vectors are aligned in each model pair by WM. Since the models $\bm{\theta}_{b}$ and $\bm{\theta}_{c}$ are matched to $\bm{\theta}_{a}$ through WM, it is expected that $R(\bm{\theta}_{a},\pi_{b}(\bm{\theta}_{b}))$ and $R(\bm{\theta}_{a},\pi_{c}(\bm{\theta}_{c}))$ would be large. On the other hand, although $\bm{\theta}_{b}$ and $\bm{\theta}_{c}$ were not explicitly aligned, $R(\pi_{b}(\bm{\theta}_{b}),\pi_{c}(\bm{\theta}_{c}))$ is clearly greater than zero, indicating that the directions of these two singular vectors are indirectly aligned by WM. From this result, the barrier between the models $\pi_{b}(\bm{\theta}_{b})$ and $\pi_{c}(\bm{\theta}_{c})$ is expected to be small. To confirm this, Table 3 shows the barriers between $\pi_{b}(\bm{\theta}_{b})$ and $\pi_{c}(\bm{\theta}_{c})$. Table 5 shows the detailed results, and Figure 17 shows the test accuracy landscape around $\bm{\theta}_{a}$, $\pi_{b}(\bm{\theta}_{b})$, and $\pi_{c}(\bm{\theta}_{c})$. As can be seen from Table 3, the barrier between $\pi_{b}(\bm{\theta}_{b})$ and $\pi_{c}(\bm{\theta}_{c})$ is smaller with WM than with STE. This means that there is a significant difference between the principles of permutations obtained by WM and STE. Figure 17 also shows that the landscape of test accuracy is flatter around the three models with WM than with STE. Therefore, WM is likely to be more advantageous, especially for merging three or more models.
@@ 7 Conclusion
This paper analyzed why linear mode connectivity (LMC) is satisfied through permutation search with weight matching (WM). First, we demonstrated that WM does not reduce the distance between the weights of two models as significantly as previously thought. We then analyzed WM using singular value decomposition (SVD) and found that WM aligns the directions of singular vectors with large singular values, which plays a crucial role in achieving LMC. Additionally, we showed that the reason LMC is established in AM is likely the same as for WM. Finally, we discussed the difference between STE and WM from the perspective of singular vectors.
Although this paper primarily analyzed WM from the perspective of individual layers (e.g., Theorem 4.2), it remains unclear why our analysis can explain the phenomenon so effectively since the actual network is multi-layered. In the future, a more comprehensive analysis that accounts for the multi-layered structure of the network will be necessary.
Ethics Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential ethical consequences of our work, none which we feel must be specifically highlighted here.
Reproducibility
The settings for reproducing the experiments are described in Appendix D. All the proofs of the theorems are given in Appendix G.
@@ Appendix A Extended Related Work
Several studies (Garipov et al., 2018; Draxler et al., 2018; Freeman & Bruna, 2017) have found that different neural network solutions can be connected by nonlinear paths with almost no increase in loss. Nagarajan & Kolter (2019) first discovered that solutions can be connected by linear paths with an almost constant loss value when training models on MNIST with the same random initial values. Later, Frankle et al. (2020) demonstrated experimentally that LMC is not always satisfied between two SGD solutions, even with the same initial parameters, depending on the datasets and model architectures. However, they also showed that if a single model is trained for a certain period and then two models are trained independently from this pre-trained model as a starting point, they are linearly mode-connected. Furthermore, Frankle et al. (2020) explored the relationship between LMC and the lottery-ticket hypothesis (Frankle & Carbin, 2019). Entezari et al. (2022) conjectured that LMC is satisfied with a high probability between two SGD solutions by accounting for permutation symmetries in the hidden layers. Subsequently, Ainsworth et al. (2023) proposed a WM method by formulating neuron alignment as a bipartite graph matching problem and solving it approximately. Later, Guerrero-Peña et al. (2023) suggested using Sinkhorn’s algorithm to solve the WM directly. Some previous studies (Ainsworth et al., 2023; Crisostomi et al., 2024) have also proposed permutation search methods for achieving LMC between multiple models. However, all of these methods reduce the $L^{2}$ distances between the models, and no methods have been proposed that use information on the loss function, such as STE. Therefore, in this paper, we created a pair of models and performed a permutation search for each pair to clarify the differences between WM and STE in a fair manner. The investigation of permutation search methods for multiple models is a future work.
@@
While several papers (Venturi et al., 2019; Nguyen et al., 2019; Nguyen, 2019; Kuditipudi et al., 2019) have discussed nonlinear mode connectivity, there is little theoretical analysis on LMC. Ferbach et al. (2024) provided an upper bound on the minimal width of the hidden layer to satisfy LMC. However, to prove this, they assumed the independence of all neuron’s weight vectors inside a given layer. It is unlikely that this assumption holds for models after training. Partially similar to our paper, Singh et al. (2024) demonstrated that the barrier value can be approximated using a second-order Taylor approximation for the case of spawning (Zhou et al., 2023). However, they have not validated this approach for permutations, and we revealed that a second-order Taylor approximation fails to accurately estimate the barrier value in the case of permutations. Zhou et al. (2023) introduced the concept of layerwise linear feature connectivity (LLFC) and showed that LLFC implies LMC. Additionally, Zhou et al. (2023) demonstrated that if weak additivity for ReLU activation and the commutativity property are satisfied, then both LLFC and LMC are satisfied. However, we show that the $L^{2}$ distance between the models after permutation is not close enough to satisfy the commutativity property. This motivated us to investigate the relationship between LMC and WM.
@@
Empirically, some previous studies (Tukan et al., 2020; Arora et al., 2018; Alvarez & Salzmann, 2017; Yu et al., 2017; Denton et al., 2014) on the compression of trained DNN models have pointed out that even when the weights are replaced with low-rank matrices, the accuracy does not decrease significantly. This suggests that SGD has an implicit bias toward reducing the model weights to low-rank. Yunis et al. (2024) explored the reduction of rank during SGD-based training and the orientation of top singular vectors via SVD on the weights. While this is relevant to our paper, Yunis et al. (2024) do not examine the effects of permutations in detail. Moreover, it does not discuss the interplay between hidden layer inputs and top singular vectors under permutations. Galanti et al. (2022) and Timor et al. (2023) further discuss the low-rank effect on the weights of trained models introduced by weight decay and a small initialization scale. In particular, Galanti et al. (2022) state that the weights become more low-rank by strengthening the weight decay and increasing the learning rate, and this is expected to help establish LMC via WM (which we experimentally confirm in Section H.5).